# An ordering for the strength of functional dependence

### v1

@@ 2.1 Decreasing rearrangements
For verifying the conditional convex order, we use the Hardy-Littlewood-Polya theorem which characterizes the standard convex order by the Schur order for functions. The latter is defined by comparing decreasing rearrangements as follows. For $d,d^{\prime}\in\mathbb{N}\,$, let $f\colon(0,1)^{d}\to\mathbb{R}$ and $g\colon(0,1)^{d^{\prime}}\to\mathbb{R}$ be integrable functions. Then the Schur order $f\prec_{S}g$ is defined by
@@
The Schur order is characterized by the following multivariate version of the Hardy-Littlewood-Polya theorem; see [19, Theorem 2.5].
Lemma 2.1 (Hardy-Littlewood-Polya theorem)
Let $f\colon(0,1)^{d}\to\mathbb{R}$ and $g\colon(0,1)^{d^{\prime}}\to\mathbb{R}$ be integrable functions. Then the following statements are equivalent:
  1. (a)
@@
To apply the Hardy-Littlewood-Polya theorem for a characterization $\preccurlyeq_{ccx}$, we use quantile transformations to get back to the unit interval/hypercube. Recall that $q_{Y}$ is the left-continuous quantile function of $Y$. We also need the multivariate quantile transform ${\bf q}_{\mathbb{X}}\colon(0,1)^{p}\to\mathbb{R}^{p}$ defined in (36). It satisfies $\mathbb{X}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}{\bf q}_{\mathbb{X}}(\mathbb{U})$ for $\mathbb{U}$ uniform on $(0,1)^{p}$. Now, define for $v\in(0,1)$ the function $\eta_{Y|\mathbb{X}}^{v}\colon(0,1)^{p}\to\mathbb{R}$ by
@@
Due to the following result, the conditional convex order can be characterized by a comparison of the functions $\eta_{Y|\mathbb{X}}^{v}$ and $\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v}$ in Schur order. We will use this characterization to verify the conditional convex order in various models; in particular, Theorem 1.8 on additive error models and Proposition 2.5 on bivariate Bernoulli distributions rely on it.
@@
    In our preliminary work [3, 5], we defined the so-called Schur order for conditional distributions by Theorem 2.2(ii) to establish ordering results for Chatterjee’s rank correlation in the setting of bivariate copula models. Here, we consider the conditional convex order, which is equivalent to the Schur order for conditional distributions but conceptually simpler.
@@
In the following example, we determine the decreasing rearrangement of $\eta_{Y|\mathbb{X}}^{v}$ in (6) for bivariate Bernoulli distributions.
@@
Consider a bivariate Bernoulli distributed random vector $(Y,X)$ taking on values $(0,0),(0,1),(1,0)$ and $(1,1)$. Denote by $p_{ji}:=P(Y=j,X=i)$, $j,i\in\{0,1\}$, the corresponding probability of occurrence such that $q:=P(Y=1)=p_{10}+p_{11}\in(0,1)$ and $p:=P(X=1)=p_{01}+p_{11}\in(0,1)$. With this notation, the bivariate Bernoulli distribution has in total $6$ parameters with $3$ degrees of freedom, where $q$ determines the distribution of $Y$ and where $\alpha:=p_{00}/(1-p)$ and $\beta:=p_{01}/p$ determine the conditional probabilities $P(Y=0\mid X=0)$ and $P(Y=0\mid X=1)$. The conditional distribution function $\eta_{Y|X}^{v}$ in (6) and its decreasing rearrangement $(\eta_{Y|X}^{v})^{\ast}$ in (9) are given by
@@
The following result characterizes the conditional convex order for bivariate Bernoulli-distributed random vectors.
@@
Let $(Y,X)$ and $(Y^{\prime},X^{\prime})$ be Bernoulli distributed random vectors with parameters as in Example 2.4. Then the following statements are equivalent:
@@
    $X$ and $Y$ are independent if and only if $\alpha=\beta$.
@@
    $Y$ perfectly depends on $X$ if and only if $\alpha\wedge\beta=0$ and $\alpha\vee\beta=1$.
@@
    $X$ and $Y$ are comonotone if and only if $\alpha=\tfrac{1-q}{1-p}\wedge 1$.
@@
    In (ii), the condition $q=q^{\prime}$ follows from the marginal constraint of the conditional convex order, i.e., $\{0,1-q,1\}=\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}=\{0,1-q^{\prime},1\}$. Large and small values of $\alpha$ and $\beta$ yield strong couplings between $Y$ and $X$, and hence a stronger functional dependence in conditional convex order than moderate values of $\alpha$ and $\beta$.
@@
    In the characterization of perfect dependence in (iv), the condition $\alpha\wedge\beta=0$ and $\alpha\vee\beta=1$ corresponds to the case $Y=X$ or Y=1XY=1-X almost surely. In these cases, it is $p=q$ and p=1qp=1-q, respectively. Due to (v) and (vi), perfect dependence of $Y$ on $X$ implies comonotonicity or countermonotonicity of $(Y,X)$; however, the converse is not true because $(Y,X)$ being comonotone (resp. countermonotone) only means that $Y=q_{Y}(U)$ and $X=q_{X}(U)$ (resp. $X=q_{X}(1-U)$) a.s. for $U$ uniform on $(0,1)$. Consequently, the extreme cases of $\leq_{c}$ and $\preccurlyeq_{ccx}$ generally differ.
@@ 2.2 A Dimension reduction principle
In this subsection, we derive a characterization of the conditional convex order in terms of the concordance order for bivariate SI distributions. To this end, we perform a dimension reduction based on decreasing rearrangements. For convenience, we work with cumulative probabilities and distribution functions instead of survival probabilities and survival functions.
To derive various properties of the conditional convex order, we denote, for all $y\in\mathbb{R}$, by $u\mapsto F_{u}(y)$ the decreasing rearrangement of $\mathbb{u}\mapsto P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))$. Recall that $u\mapsto F_{u}(y)$ is $\lambda$-almost surely uniquely determined. In the sequel, we will consider versions such that $y\mapsto F_{u}(y)$ is a distribution function for all $u\in(0,1)$; see Lemma B.1 for the existence of such rearrangements. Setting $y=q_{Y}(v)$ for $v\in(0,1)$, the function $u\mapsto F_{u}(q_{Y}(v))$ is the decreasing rearrangement of $\eta_{Y|\mathbb{X}}^{v}$ in (6), i.e.
$(\eta_{Y|\mathbb{X}}^{v})^{*}(u)=F_{u}(q_{Y}(v))\quad\text{for all }u\in[0,1].$ (9)
@@
$q_{Y|\mathbb{X}}^{\uparrow u}(t):=F_{u}^{-1}(t),\quad t\in(0,1],$ (10)
which we refer to as rearranged quantile transform. To explain the choice of the notation in (10), let $V$ be uniform on $(0,1)$. First, note that $q_{Y|\mathbb{X}}^{\uparrow u}(V)$ is a random variable with distribution function $F_{u}$. Since, by construction, $F_{u}(y)$ is decreasing in $u$ for all $y\in\mathbb{R}$, the family $\{q_{Y|\mathbb{X}}^{\uparrow u}(V)\}_{u\in[0,1]}$ of random variables induced by (10) is increasing in $u$ with respect to the stochastic order. This means that $u_{1}\leq u_{2}$ implies $\mathbb{E}f(q_{Y|\mathbb{X}}^{\uparrow u_{1}}(V))\leq\mathbb{E}f(q_{Y|\mathbb{X}}^{\uparrow u_{2}}(V))$ for all increasing functions $f$ such that the expectations exist.
@@
Figure 2: Example for the construction of the bivariate SI random vector $(q_{Y|X}^{\uparrow U}(V),U)$ in Proposition 2.7(ii) when a discrete bivariate random vector $(Y,X)$ mapping into a grid $\{a_{1},\ldots,a_{4}\}\times\{b_{1},\ldots,b_{4}\}$ is given: The top left matrix illustrates the mass distribution of $(Y,X)$; the bottom left matrix describes the associated conditional distribution functions $\eta_{Y|X}^{v}\,;$ the bottom right matrix defines the conditional distribution function $v\mapsto F_{u}(q_{Y}(v))$ that is defined in (9) and decreasing in $u$; the top right matrix describes the mass distribution of the SI random vector $(q_{Y|X}^{\uparrow U}(V),U)$ defined by (10).
Using the rearranged quantile transform constructed above, the following result shows that every random vector $(Y,\mathbb{X})$ can be reduced to a bivariate SI random vector having the same strength of functional dependence in the sense of the conditional convex order.
Proposition 2.7 (Dimension reduction to bivariate SI distributions)
@@
    $q_{Y|\mathbb{X}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y\,$,
@@
    $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ is a bivariate SI random vector,
@@
    $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=_{ccx}(Y,\mathbb{X})\,.$
@@
  1. (a)
    Due to Proposition 2.7, the rearranged quantile transform relates a $(1+p)$-dimensional random vector $(Y,\mathbb{X})$ to a bivariate SI random vector such that the dependence information in the sense of the conditional convex order remains invariant. Consequently, the strength of functional dependence is preserved in the sense that all dependence measures which are monotone in $\preccurlyeq_{ccx}$ are invariant under this transformation. We refer to Section 4 for classes of $\preccurlyeq_{ccx}$-increasing dependence measures. Our transformation extends the rearrangement of bivariate copulas in [4, Proposition 3.17] and [62, Theorem 2.3] to arbitrary marginal distributions and to a multivariate random vector $\mathbb{X}$.
  2. (b)
    For $A={\overline{\mathsf{Ran}(F_{Y})}}$, $(Y,\mathbb{X})$ and $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ are minimal elements in $\mathcal{R}_{A}$ if and only if $Y$ and $\mathbb{X}$ are independent or, equivalently, if $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ and $U$ are independent. Similarly, they are maximal elements in $\mathcal{R}_{A}$ if and only if $Y$ perfectly depends on $\mathbb{X}$, which is equivalent to $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ being perfectly dependent on $U$. Since $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ is SI in $U$, the latter property is equivalent to comonotonicity of $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ and $U$.
For a bivariate random vector with finite support, the construction of the rearranged quantile transform in (10) is explained in Figure 2. For a multivariate normal random vector, we illustrate the dimension reduction principle and the construction of the bivariate SI random vector $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ in the following example. Note that we use the basic ideas in this example to study sufficient conditions for verifying the conditional convex order in Section 3.2.
@@
Let $(Y,\mathbb{X})\sim N(\mathbf{0},\Sigma)$ be a $(1+p)$-dimensional normal random vector. In the sequel, we construct the bivariate SI random vector $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ in Proposition 2.7 based on the dimenion reduction principle in (9) and (10). Consider the decomposition of the covariance matrix into
@@
Define $S:=A\mathbb{X}$ for $A:=\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}/\sqrt{\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}}\,$, where $\Sigma_{\mathbb{X}}^{-}$ denotes a generalized inverse of $\Sigma_{\mathbb{X}}$ such as the Moore-Penrose inverse. Key for the dimension reduction principle and for the construction of the rearranged quantile transform $q_{Y|\mathbb{X}}^{\uparrow u}$ in (10) is the identity
@@
see [2, Proof of Proposition 2.7]. For $k:=\mathsf{rank}(\Sigma_{\mathbb{X}})\,$, let $B$ be a $p\times k$ matrix such that $\Sigma_{\mathbb{X}}=BB^{T}.$ Then the transformed conditional distribution function $\eta_{Y|\mathbb{X}}^{v}$ in (6) is given by
$\begin{split}\eta_{Y|\mathbb{X}}^{v}(\mathbb{u})&=P(Y\leq\sigma_{Y}\Phi^{-1}(v)\mid\mathbb{X}=B{\bf\Phi}^{-1}(\mathbb{u}))\\ &=P(Y\leq\sigma_{Y}\Phi^{-1}(v)\mid S=AB{\bf\Phi}^{-1}(\mathbb{u}))\\ &=P(Y\leq\sigma_{Y}\Phi^{-1}(v)\mid S=\Phi^{-1}(u))\end{split}$ (13)
@@
which implies that $Y$ is SI in S;S\,; see e.g. [50]. Let $\mathbb{U}$ be a random vector that is uniform on $(0,1)^{k}\,$, and define $U:=\Phi(S)=\Phi(AB{\bf\Phi}^{-1}(\mathbb{U}))$. Then $U$ is uniform on $(0,1)$ because $S$ is standard normal. The decreasing rearrangement of $\eta_{Y|\mathbb{X}}^{v}$ in (13) is given by
@@
for all $z\in\mathbb{R}\,.$ The second equality in (16) follows from (13). For the first equality in (16), we use that $U$ is uniform on (0,1);(0,1)\,; for the third equality we use that $\mathbb{U}$ is uniform on $(0,1)^{k}\,.$ For fixed $u\in(0,1)\,$, the rearranged distribution function $F_{u}$ in (9) is given by
@@
$q_{Y|\mathbb{X}}^{\uparrow u}(t)=F_{u}^{-1}(t)=F_{Y|S=\Phi^{-1}(u)}^{-1}(t)\,,\quad t\in(0,1)\,.$ (18)
Finally, the bivariate SI random vector in Proposition 2.7(ii) is given by
$(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=(F_{Y|S=\Phi^{-1}(U)}^{-1}(V),U).$ (19)
Note that the left-hand side in (19) satisfies $(q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=_{ccx}(Y,\mathbb{X})$, which follows from (13) and (17) and verifies Proposition 2.7(iii). Due to (12), we know that $(Y,\mathbb{X})=_{ccx}(Y,S)$. Hence, the reduced random vectors in Theorem 1.6 satisfy
$(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y}\circ Y,\Phi(S))$ (20)

### v2

@@ 2.1 Decreasing rearrangements
In this subsection, we characterize the conditional convex order through rearrangements. This concept plays a key role for verifying $\preccurlyeq_{ccx}$ in several settings. Large parts of the paper, in particular, Sections 2.2, 3, and 4.3, as well as Theorem 1.6iii and Theorem 1.8, are based on rearrangements.
For $d,d^{\prime}\in\mathbb{N}\,$, let $f\colon(0,1)^{d}\to\mathbb{R}$ and $g\colon(0,1)^{d^{\prime}}\to\mathbb{R}$ be integrable functions. Then the Schur order $f\prec_{S}g$ is defined by
@@
The Schur order is characterized by the following multivariate Hardy-Littlewood-Polya theorem; see [19, Theorem 2.5].
Lemma 2.1 (Hardy–Littlewood–Polya theorem)
The following are equivalent:
  1. (i)
@@
To apply the Hardy-Littlewood-Polya theorem for characterizing $\preccurlyeq_{ccx}$, we use quantile transformations to get back to the unit interval/hypercube. Recall that $q_{Y}$ is the left-continuous quantile function of $Y$. We also need the multivariate quantile transform ${\bf q}_{\mathbb{X}}\colon(0,1)^{p}\to\mathbb{R}^{p}$ defined in (A.1). It satisfies $\mathbb{X}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}{\bf q}_{\mathbb{X}}(\mathbb{U})$ for $\mathbb{U}$ uniform on $(0,1)^{p}$. Now, define for $v\in(0,1)$ the function $\eta_{Y|\mathbb{X}}^{v}\colon(0,1)^{p}\to\mathbb{R}$ by
@@
Due to the following result, the conditional convex order can be characterized by a comparison of the functions $\eta_{Y|\mathbb{X}}^{v}$ and $\eta_{Y^{\prime}|\mathbb{X}^{\prime}}^{v}$ in Schur order.
@@
    In our preliminary work [5, 4], we introduced the so-called Schur order for conditional distributions through the characterization in Theorem 2.2ii and derived ordering results for Chatterjee’s rank correlation within bivariate copula models. In this paper, we consider the conditional convex order, which is equivalent to the Schur order for conditional distributions but conceptually simpler.
@@
To give a first illustration how to verify $\preccurlyeq_{ccx}$ by Theorem 2.2, we construct the decreasing rearrangement of $\eta_{Y|\mathbb{X}}^{v}$ in (9) for bivariate Bernoulli distributions.
@@
Consider a bivariate Bernoulli distributed random vector $(Y,X)$ taking on values $(0,0),(0,1),(1,0)$ and $(1,1)$. Denote by $p_{ji}:=P(Y=j,X=i)$, $j,i\in\{0,1\}$, the corresponding probability of occurrence such that $q:=P(Y=1)=p_{10}+p_{11}\in(0,1)$ and $p:=P(X=1)=p_{01}+p_{11}\in(0,1)$. With this notation, the bivariate Bernoulli distribution has in total $6$ parameters with $3$ degrees of freedom, where $q$ determines the distribution of $Y$ and where $\alpha:=p_{00}/(1-p)$ and $\beta:=p_{01}/p$ determine the conditional probabilities $P(Y=0\mid X=0)$ and $P(Y=0\mid X=1)$. The conditional distribution function $\eta_{Y|X}^{v}$ in (9) and its decreasing rearrangement $(\eta_{Y|X}^{v})^{\ast}$ are given by
@@
To characterize $\preccurlyeq_{ccx}$ for bivariate Bernoulli distributions in terms of Theorem 2.2, we need to compare integrals of the decreasing rearrangements in (11); see Figure 1. The following result further illustrates the different concepts of monotone and functional dependence.
@@
Let $(Y,X)$ and $(Y^{\prime},X^{\prime})$ be Bernoulli distributed random vectors with parameters as in Example 2.4. Then the following two statements are equivalent:
@@
    $X$ and $Y$ are independent if and only if $\alpha=\beta$;
@@
    $Y$ perfectly depends on $X$ if and only if $\alpha\wedge\beta=0$ and $\alpha\vee\beta=1$;
@@
    $X$ and $Y$ are comonotone if and only if $\alpha=\tfrac{1-q}{1-p}\wedge 1$;
@@
    In ii, the condition $q=q^{\prime}$ follows from the marginal constraint of the conditional convex order, i.e., $\{0,1-q,1\}=\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}})}=\{0,1-q^{\prime},1\}$. Large and small values of $\alpha$ and $\beta$ yield strong couplings between $Y$ and $X$, and hence a stronger functional dependence in conditional convex order than moderate values of $\alpha$ and $\beta$.
@@
    In the characterization of perfect dependence in iv, the condition $\alpha\wedge\beta=0$ and $\alpha\vee\beta=1$ corresponds to the case $Y=X$ or Y=1XY=1-X almost surely. In these cases, it is $p=q$ and p=1qp=1-q, respectively. Due to v and vi, perfect dependence of $Y$ on $X$ implies comonotonicity or countermonotonicity of $(Y,X)$; however, the converse is not true because $(Y,X)$ being comonotone (resp. countermonotone) only means that $Y=q_{Y}(U)$ and $X=q_{X}(U)$ (resp. $X=q_{X}(1-U)$) a.s. for $U\sim U(0,1)$. Consequently, the extreme cases of $\leq_{c}$ and $\preccurlyeq_{ccx}$ generally differ; see also Section 5.1.
@@ 2.2 A dimension reduction principle
In this subsection, we reduce $(Y,\mathbb{X})$ to a $\preccurlyeq_{ccx}$-equivalent bivariate SI random vector using the concept of decreasing rearrangements. The dimension reduction principle underlies Theorem 1.6iii and Theorem 1.8 as well as the results in Sections 3 and 4.3. For convenience, we work with cumulative probabilities and distribution functions instead of survival probabilities and survival functions.
To transform $(Y,\mathbb{X})$ into a $\preccurlyeq_{ccx}$-equivalent bivariate random vector, we denote by $0,1]\ni u\mapsto F_{u}(y)$ the decreasing rearrangement of $\mathbb{u}\mapsto P(Y\leq y\mid\mathbb{X}=q_{\mathbb{X}}(\mathbb{u}))$. Recall that $u\mapsto F_{u}(y)$ is $\lambda$-almost surely uniquely determined. In the sequel, we will consider versions such that $y\mapsto F_{u}(y)$ is a distribution function for all $u\in(0,1)$; see Lemma B.1 for the existence of such rearrangements. Setting $y=q_{Y}(v)$ for $v\in(0,1)$, we observe that the function $u\mapsto F_{u}(q_{Y}(v))$ is the decreasing rearrangement of $\eta_{Y|\mathbb{X}}^{v}$ in (9), i.e.
$(\eta_{Y|\mathbb{X}}^{v})^{*}(u)=F_{u}(q_{Y}(v))\quad\text{for all }u\in(0,1).$ (12)
@@
$q_{Y;\mathbb{X}}^{\uparrow u}(t):=F_{u}^{-1}(t),\quad t\in(0,1),$ (13)
which we refer to as rearranged quantile transform. To explain the choice of the notation in (13), let $V\sim U(0,1)$. First, note that $q_{Y;\mathbb{X}}^{\uparrow u}(V)$ is a random variable with distribution function $F_{u}$. Since, by construction, $F_{u}(y)$ is decreasing in $u$ for all $y\in\mathbb{R}$, the family $\{q_{Y;\mathbb{X}}^{\uparrow u}(V)\}_{u\in[0,1]}$ of random variables is increasing in $u$ with respect to the stochastic order. This means that $u_{1}\leq u_{2}$ implies $\mathbb{E}f(q_{Y;\mathbb{X}}^{\uparrow u_{1}}(V))\leq\mathbb{E}f(q_{Y;\mathbb{X}}^{\uparrow u_{2}}(V))$ for all increasing functions $f$ such that the expectations exist.
@@
Figure 2: Example for the construction of the bivariate SI random vector $(q_{Y;X}^{\uparrow U}(V),U)$ in Theorem 2.7ii when a discrete bivariate random vector $(Y,X)$ maps into a grid $\{a_{1},\ldots,a_{4}\}\times\{b_{1},\ldots,b_{4}\}$: The top left matrix illustrates the mass distribution of $(Y,X)$; the bottom left matrix describes the associated conditional distribution functions $\eta_{Y|X}^{v}\,;$ the bottom right matrix is the rearranged conditional distribution function $v\mapsto F_{u}(q_{Y}(v))$ which is defined in (12) and decreases in $u$; the top right matrix describes the mass distribution of the SI random vector $(q_{Y;X}^{\uparrow U}(V),U)$ defined by (13).
Using the rearranged quantile transform constructed above, the following result shows that every random vector $(Y,\mathbb{X})$ can be reduced to a $\preccurlyeq_{ccx}$-equivalent bivariate SI random vector.
Theorem 2.7 (Dimension reduction to bivariate SI distributions)
@@
    $q_{Y;\mathbb{X}}^{\uparrow U}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y\,$,
@@
    $(q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ is a bivariate SI random vector,
@@
    $(q_{Y;\mathbb{X}}^{\uparrow U}(V),U)=_{ccx}(Y,\mathbb{X})\,.$
@@
Due to Theorem 2.7, the rearranged quantile transform relates a $(1+p)$-dimensional random vector $(Y,\mathbb{X})$ to a bivariate SI random vector such that the dependence information in the sense of $\preccurlyeq_{ccx}$ remains invariant. Consequently, the strength of functional dependence is preserved in the sense that all dependence measures which are monotone in $\preccurlyeq_{ccx}$ are invariant under this transformation. The reduced random vector in (7) is based on this dimension-reduction principle. Note that our transformation extends the rearrangement of bivariate copulas in [2, Proposition 3.17] and [62, Theorem 2.3] to arbitrary marginal distributions and to any multivariate random vector $\mathbb{X}$.
For a bivariate random vector with finite support, the construction of the rearranged quantile transform in (13) is explained in Figure 2. For a multivariate normal random vector, we perform the dimension reduction and the construction of the bivariate SI random vector $(q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ in the following example. Note that we use the basic ideas from this example to study in Section 3.2 sufficient conditions for verifying $\preccurlyeq_{ccx}$.
@@
Let $(Y,\mathbb{X})\sim N(\mathbf{0},\Sigma)$ be a $(1+p)$-dimensional normal random vector. In the sequel, we construct the bivariate SI random vector $(q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ in Theorem 2.7 based on the dimenion reduction principle in (12) and (13). Consider the decomposition of the covariance matrix into
@@
Define $S:=A\mathbb{X}$ for $A:=\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}/\sqrt{\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}}\,$, where $\Sigma_{\mathbb{X}}^{-}$ denotes a generalized inverse of $\Sigma_{\mathbb{X}}$ such as the Moore-Penrose inverse. Key for the dimension reduction principle and for the construction of the rearranged quantile transform $q_{Y;\mathbb{X}}^{\uparrow u}$ in (13) is the identity
@@
see [3, Proof of Proposition 2.7]. For $k:=\mathsf{rank}(\Sigma_{\mathbb{X}})\,$, let $B$ be a $p\times k$ matrix such that $\Sigma_{\mathbb{X}}=BB^{T}.$ Then the transformed conditional distribution function $\eta_{Y|\mathbb{X}}^{v}$ in (9) is given by
$\begin{split}\eta_{Y|\mathbb{X}}^{v}(\mathbb{u})&=P(Y\leq q_{Y}(v)\mid\mathbb{X}=B{\bf\Phi}^{-1}(\mathbb{u}))\\ &=P(Y\leq\sigma_{Y}\Phi^{-1}(v)\mid S=AB{\bf\Phi}^{-1}(\mathbb{u}))\\ &=P(Y\leq\sigma_{Y}\Phi^{-1}(v)\mid S=\Phi^{-1}(u))\end{split}$ (16)
@@
which implies that $Y$ is SI in S;S\,; see e.g. [52]. Let $\mathbb{U}$ be a random vector that is uniform on $(0,1)^{k}$ and define $U:=\Phi(S)=\Phi(AB{\bf\Phi}^{-1}(\mathbb{U}))$. Then $U$ is uniform on $(0,1)$ because $S$ is standard normal. The decreasing rearrangement of $\eta_{Y|\mathbb{X}}^{v}$ in (16) is given by
@@
for all $z\in\mathbb{R}\,.$ The second equality in (19) follows from (16). For the first equality in (19), we use that $U$ is uniform on $(0,1)$; for the third equality we use that $\mathbb{U}$ is uniform on $(0,1)^{k}\,.$ For fixed $u\in(0,1)\,$, the rearranged distribution function $F_{u}$ in (12) is given by
@@
$q_{Y;\mathbb{X}}^{\uparrow u}(t)=F_{u}^{-1}(t)=F_{Y|S=\Phi^{-1}(u)}^{-1}(t)\,,\quad t\in(0,1)\,.$ (21)
Finally, the bivariate SI random vector in Theorem 2.7ii is given by
$(q_{Y;\mathbb{X}}^{\uparrow U}(V),U)=(F_{Y|S=\Phi^{-1}(U)}^{-1}(V),U).$ (22)
Note that the left-hand side in (22) satisfies $(q_{Y;\mathbb{X}}^{\uparrow U}(V),U)=_{ccx}(Y,\mathbb{X})$, which follows from (16) and (20) and verifies Theorem 2.7iii. Due to (15), we know that $(Y,\mathbb{X})=_{ccx}(Y,S)$. Hence, the reduced random vectors in Theorem 1.6 iii satisfy
$(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y}\circ Y,\Phi(S))$ (23)